OpenStudy (aaronandyson):

The sum of a number and its reciprocal is 4.25. Find the number. @loser66 @sepeario

OpenStudy (sepeario):

\[4.25 = x+ \frac{ 1 }{ x }\]

OpenStudy (sepeario):

Are you possibly learning the quadratic formula right now?

OpenStudy (sepeario):

Because all of these questions involve the quadratic formula.

OpenStudy (sepeario):

\[4.25x=x^2+1\] and then \[4.25x+x^2-1=0\]

OpenStudy (aaronandyson):

No. We can only use the formula when the question "SAYS" round off your answer.

OpenStudy (sepeario):

There is pretty much no other way of working it out other than guess and check .

OpenStudy (aaronandyson):

I can't use it as the question is not asking me to.

OpenStudy (aaronandyson):

-1/4 and 4?